Schedule September 2008
<!FONT SIZE=-1> return to main 1401 Restoration Page
go to Team Bios
Ed took a vacation from posting weekly details -
Wed September 24 - general, Thurs September 25 - tape team
Wed September 24 - generalPresent were: Ron Williams, Bob Erickson, Frank King, Allen Palmer, Glenn Lea, Joe Preston, Stan Paddock, Judith Haemmerle, Ed Thelen, Robert Garner
Following report from Robert Garner
- DE 1401
- Fully functional and 100% operational, for several consecutive months now! (Not to be upstaged by the new-kid CT 1401. ;-)
- Allen noted defective power supply on 2nd Model V 729. 3V instead of 12V. Ed investigating.
- CT 1401
- - Most instructions operational in 1401 CPU, including multiply. (no bad SMS cards found for most CPU instructions).
- - 1402 card reading is generating read check parity compare errors on 12'th row holes.
- - Ron & Robert can trigger scope on failures, but after several sessions, they still haven't found root cause.
- - Occasionally reading a card results in a destructive storage scan (frustrating).
- - Joe and Stan discovered Model II 729 burnt contacts in vacuum column bellow switches.
- - Frank noted issues in 1403 printer high-speed roller feed shaft (pictured)
Thurs September 25 - tape teamFrom Bob Feretich
1401 TAU debug status 9/25/08 (Jeff & I)
We replaced 2 burnt out indicator lamps and 3 bad cards.
- We are now able to write records to tape from the TAU CE panel. The data was received correctly by the PC. The TAU was happy with the wrap read back of the write data. No errors are reported.
- The Rewind switch also seems to work.
- The Rewind & Unload switch does not work.
Sat. September 27 - generalAttendees: Ron Williams, Bob Erickson
- Worked on 1402 read check problem in 1401 logic
- Bob noted that the 1402, after decades of use, needs a thorough cleaning and rehab. (As soon as CT system is up, and since the DE 1401 is operational, we should consider a full mechanical overhaul/refurb of the CT 1402!)
Go to October 2008
go to Team Bios